BCAL Reconstruction Meeting 2015-01-22

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m. EDT

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204


  1. BCAL Reconstruction Issues
  2. BCAL Reconstruction Algorithms
  3. Will's reconstructed energy info

Action Items

  1. Feedback to Elton - GlueX-doc-2618 and GlueX-doc-2622 (All)
  2. 12 bit ADC - ask Sean Dobbs (Zisis)
  3. Elton's low-level coefficients into ccdb - ask Sean Dobbs (Zisis)
  4. Match with tracks - ask Paul, Simon (Zisis)
  5. Check new dark hits code after David's adjustments (Tegan)
  6. Shower curvature tables: compare KLOE to IU codes: check gap in cluster handling between codes (Tegan, Andrei)
  7. z-coordinate determination from up/down amplitude ratio (Tegan, Andrei)
  8. Cleanup: (Tegan, Will L.)
    1. Set variable “failures” to a number other than zero (e.g. -9999 or similar)
    2. Comb code for constants to be pulled out into a database
  9. Measure data pedestals using LV=0 setting (Mark)
  10. Sampling fraction tables (Tegan)


  1. Announcements
  2. Action Items
  3. Prioritize BCAL Reconstruction Issues
  4. Commissioning
    1. Pedestal skims (Tegan)
  5. Calibration
  6. Simulations
    1. BDT effort (Tegan/Justin)
  7. Any other business


Attendees: David, Will M. (JLab), Will L. (CMU), Justin (MIT), Christine, George (Athens), Tegan, Ahmed, Andrei, Zisis (Regina)

  1. Announcements: Justin was at Regina last week working with the group on BDT. An undergraduate student (Laura Teigrob) will be working with Tegan on the pedestals, a.o.
  2. Action Items:
    1. Elton requested feedback on his two documents.
    2. No reply from Sean yet.
    3. Sean and Mark will work together on this; a) single FADC time slope constant, b) t-w corrections will follow a four-parameter equation, c) postpone TDC work on BCAL for now.
    4. Track matching: Justin reported that this is going well and provide some info: Run 2400. Put your mouse over the Matching2 box and you'll see the BCAL DeltaPhi and DeltaZ matches with tracks. Or go to the direct link Histo Matching
    5. Items 5-10 no progress.
  3. Prioritize BCAL Reconstruction Issues: Skip, since Elton and Mark are not present.
  4. Commissioning
    1. Pedestal skims (Tegan): Tegan provided scripts to Will M. to process all runs, so that we can have a text file with the runs and no of pedestal events on the BCAL.
    2. Analysis (Will M.): energy calibration work in progress, nothing to report.
  5. Calibration
    1. Two Regina grad students will be at JLab in February to work on cosmics runs, etc: Noemi Ochoa (Feb 1 - Mar 4), Ahmed Foda (Feb 17 - Mar 4).
    2. Tests are being mounted at UofR to investigate issues connected with the reflections we see in the data.
  6. Simulations
    1. Cosmic runs with low threshold. Irina/Andrei will rerun the cosmics MC for low thresholds, in the hope of capturing the 'missing shoulders' in the energy deposition in some cells and come up with a better fit for the calibration constants. Also, the non-constant attenuation length (450 cm and up) could be incorporated, perhaps up to 550 cm. Our Working Group will need to decide what to do with the 550-750 cm range from Mark's fits.
    2. BDT effort (Tegan/Justin)
    3. Tegan and Justin reported that the info from the truth showers is now incorporated in the BDT, to examine in detail the 'bad clusters' fooling the BDT to date. The PhotonDiscim plugin is now using the truth info instead of the thrown info. The idea is also to grab the PID of anything spatially matched to a shower and put into an array for further work. There was an issue with the shower timing (picket fence histo) which Justin explained was due to a change by Richard Jones as we have no vertex info, and this was calculated from other event parameters. This affects all simulations. Justin said that it works for the full event now, but not for the BCAL showers yet.
  7. Any other business
    1. George is writing exams at U of Athens until the end of February, after which he will continue on research, data analysis and reconstruction.